OpenStudy (steve816):

I don't understand this :( Suppose (-3, 5) is a point on the graph of y = g(x) What point is on the graph of y = -3g(x-4) + 3 What point is on the graph of y = g(3x+9)

OpenStudy (steve816):

@imqwerty

OpenStudy (anonymous):

Well the line has a horizontal shift to the right 4 units so this says (-3,5) turns to (1,5)

OpenStudy (anonymous):

then has slope of -3

OpenStudy (anonymous):

which becomes (-3,-15) then shifts up 3 units upward, (-3,-12)

OpenStudy (anonymous):

i might be wrong

imqwerty (imqwerty):

g(x)=y and u got- y = -3g(x-4) + 3 y-3=-3g(x-4) (y-3)/(-3) =g(x-4) compare this with g(x)=y we know that in this^ when x=-3 y=5 so compare g(x)=y with (y-3)/(-3) =g(x-4) x-4=3 x=7 (y-3)/(-3) =5 y-3=-15 y=-12 so (7,-12) do the same with the second one :)

OpenStudy (anonymous):

it should be (-3,12) your step with x-4=3 should actually be x-4=-3